OpenStudy (anonymous):

Change the Cartesian integral int[0 to 9]int[0 to y]x dx dy into an equivalent polar integral.

OpenStudy (anonymous):

\[\int\limits_{0}^{9}\int\limits_{0}^{y}x \ dx dy\] I know the first part is \[\int\limits_{0}^{9/\sin }\] but I don't know the second part

OpenStudy (anonymous):

Yeah, this is just sort of a silly integral to transform.

OpenStudy (anonymous):

any thoughts? I have the first part: \[\int\limits_{?}^{?}\int\limits_{0}^{9/\sin (\theta)}r ^{2}\cos \theta \ dr d \theta \]

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

y=x isn't it?

OpenStudy (anonymous):

I mean for the original integral

OpenStudy (anonymous):

yes 0<x<y

OpenStudy (anonymous):

I would say you integrate from 45% until 90% so \( \large \frac{\pi}{4}< \theta< \frac{\pi}{2} \)

OpenStudy (anonymous):